# Beyond Direct Generation: A Decomposed Approach to Well-Crafted Screenwriting with LLMs

### v1

@@ Abstract
The screenplay serves as the foundation for television production, defining narrative structure, character development, and dialogue. While Large Language Models (LLMs) offer great potential in this creative process, direct end-to-end generation approaches often fail to produce well-crafted screenplays. We argue this failure stems from forcing a single model to simultaneously master two disparate capabilities: creative narrative construction and rigid format adherence. The resulting outputs may mimic superficial style but lack the deep structural integrity and storytelling substance required for professional use. To enable LLMs to generate high-quality screenplays, we introduce a decomposed framework termed Dual-Stage Refinement (DSR). Rather than relying on direct end-to-end generation, DSR explicitly decouples creative narrative generation from format conversion. In the first stage, the framework focuses on creative narrative development by transforming a brief outline into rich, novel-style prose. The second stage then refines the narrative prose into a professionally formatted screenplay, concentrating exclusively on stylistic conversion. Such separation enables the model to specialize in one distinct capability at each stage. A significant challenge in implementing DSR is the scarcity of paired outline-to-novel data required for training the first stage. We address this through a hybrid data synthesis strategy. The process begins with reverse synthesis, deconstructing existing screenplays into structured inputs (such as outlines and character profiles). Forward synthesis then leverages these inputs to generate high-quality, narratively rich novel-style texts as training targets. Extensive experiments show that in blind evaluations by professional screenwriters, screenplays generated by DSR framework achieve a $75\%$ win rate against strong baselines like Gemini-2.5-Pro and reach $82.7\%$ of human-level performance. Our work demonstrates that a decomposed generation architecture, enabled by a tailored hybrid data synthesis strategy, is a highly effective approach for specializing LLMs in complex creative domains.
@@
Hang Lei1thanks: Corresponding author.   Shengyi Zong1  Zhaoyan Li1   Ziren Zhou1,2   Hao Liu1 1Alibaba Group,  2Peking University {leihang.lh,zongshengyi.zsy,lzy434483,zhouziren.zzr,lh414475}@alibaba-inc.com
@@ 1 Introduction
The advent of LLMs is bringing new potential to screenplay generation [^0] The terms ‘screenplay’ and ‘script’ in this work refer to scripts for episodic television, not feature films.. LLMs can serve as creative assistants by handling time-consuming tasks such as generating plot variations and character biographies, thereby allowing writers to focus on higher-level creative decisions including refining story themes and developing character depth. Current LLMs are general-purpose text generators trained primarily on web corpora and novels. For practical screenwriting applications, such models must be adapted to understand screenplay-specific narrative structures, maintain story coherence across long sequences, and manage complex character development. The research challenge is therefore developing specialized AI tools that meet these requirements and enable writers to produce higher-quality screenplays more efficiently.
@@
A screenplay differs fundamentally from a novel in that it serves as a structured framework for visual storytelling. Rather than relying on detailed prose descriptions or internal monologue, screenplays employ the principle of "showing, not telling" through precise audiovisual language. This is achieved using a strict format comprising scene headings, concise action lines, and authentic dialogue to convey plot, character, and emotion. Consequently, an effective screenplay generation model must satisfy several fundamental requirements: maintaining narrative coherence, ensuring character consistency, and adhering to strict formatting conventions that distinguish screenplays from other narrative forms.
Current research in controllable text generation has made notable progress, primarily in general story writing, focusing on two main directions: enhancing controllability and incorporating commonsense knowledge. For controllability, researchers have developed models that generate stories following plot outlines (Rashkin et al., 2020) or fine-grained personalized guidance (Wang et al., 2022b). To enhance logical coherence, others have incorporated commonsense knowledge through inference (Wang et al., 2022a) or knowledge retrieval (Qin and Zhao, 2022). In the more specialized domain of screenwriting, systems such as Dramatron (Mirowski et al., 2022) have employed hierarchical prompt chaining to generate plot structures.
@@
While these prior works have advanced controllable story generation and plot construction, our work addresses a more complex challenge that requires not only narrative content and logical coherence but also audiovisual presentation. Specifically, we focus on generating production-ready screenplays that not only tell compelling stories but also present them through professional audiovisual language and formatting conventions. This requires a model to simultaneously master two disparate skills—creative narrative generation ("what to write") and rigid format conversion ("how to write"). Handling both skills jointly dramatically increases training complexity and often leads to suboptimal performance, producing outputs that superficially mimic screenplay style but lack the conciseness, visual language, and structural integrity required for professional use. This suggests that a more effective strategy would be to decouple these two skills, allowing the model to focus on one well-defined task at a time.
@@
Motivated by the above analysis, this paper introduces a decomposed framework we term Dual Stage Refinement (DSR) that breaks down screenplay generation into two distinct stages. The first stage, creative narrative development, focuses on translating a high-level outline into rich, novel-style prose, allowing the model to concentrate purely on storytelling. The second stage, stylistic format conversion, then refines this narrative prose into a professionally formatted screenplay. Such separation allows the model to master one distinct skill at each step.
@@
Implementing the two-stage framework, however, presents a significant data challenge: the lack of paired outline to novel-style prose data required to train the first stage. To overcome this hurdle, we developed a hybrid data synthesis strategy that combines reverse synthesis and forward synthesis. Reverse synthesis deconstructs existing screenplays into structured inputs such as outlines and character profiles. Forward synthesis then uses the deconstructed inputs to generate high-quality, narratively rich novel-style targets for training. This tailored data creation process enables our DSR framework to be effectively trained.
@@
To validate our framework, we conducted blind evaluations where professional screenwriters assessed screenplays on outline adherence, narrative coherence, and character depth. The results demonstrate a decisive preference for screenplays generated by our DSR framework: a $75\%$ win rate in pairwise comparisons against SOTA models like Gemini-2.5-Pro and Claude-Sonnet-4, with overall quality reaching $82.7\%$ of human-level performance, highlighting practical viability.
In summary, the main contributions of this work can be briefly summarized as follows: (1) A Decomposed Generation Framework (DSR): We propose the Dual-Stage Refinement (DSR) framework, a novel architecture that decouples the complex screenwriting task into two distinct stages: creative narrative generation and stylistic format conversion. This decomposition effectively addresses the fundamental limitations of direct end-to-end generation approaches. (2) A Scalable Hybrid Data Synthesis Strategy: We introduce a robust and scalable method to programmatically generate the large-scale, specialized training data tailored to the DSR framework. By combining reverse engineering and forward synthesis techniques, this strategy effectively resolves the critical data scarcity bottleneck inherent to screenwriting tasks. (3) Comprehensive Validation: Through extensive experiments, including blind evaluations conducted by professional screenwriters, we demonstrate that our specialized DSR model significantly outperforms strong generalist baselines. These results validate the effectiveness of our decomposed architecture for structured creative generation tasks.
@@ 2 Related Works
Screenplay Generation. The automated generation of screenplays, a task requiring both creative narrative construction and strict structural adherence, has attracted increasing attention in recent years. Early approaches before modern LLMs predominantly relied on retrieval-based methods. For instance, Zhu et al. (2022) addressed guideline-based script generation by selecting dialogue lines from a candidate pool, introducing the GraphMovie dataset to facilitate this task.
With the advent of LLMs, research has shifted toward hierarchical, human-in-the-loop frameworks. Dramatron (Mirowski et al., 2022) exemplifies this paradigm by enabling interactive co-writing: it first generates high-level structural elements such as loglines before producing dialogue. Similarly, IBSEN (Han et al., 2024) employs a multi-agent "director-actor" framework to generate drama scripts under human supervision. While innovative, these approaches position LLMs primarily as collaborative assistants rather than autonomous creators, requiring substantial user involvement to ensure quality and coherence.
Crucially, independent evaluations of LLM narrative capabilities reveal why such supervision remains necessary. A notable study by (Tian et al., 2024) found that while models like GPT-4 excel at surface-level qualities (e.g., grammar, stylistic consistency), they consistently fail at deeper "narrative intelligence," producing clichéd plots and flat characters in end-to-end generation scenarios. This finding confirms our observation: the poor performance of end-to-end models arises from the inherent difficulty of simultaneously managing creative narrative construction and rigid structural constraints. Unlike collaborative frameworks, our work directly tackles this core generation challenge to achieve autonomous, high-quality screenplay production.
Planning-based Narrative Generation. Beyond end-to-end generation, researchers have explored two major directions: planning-based frameworks and iterative refinement methods.
The "plan-and-write" approach introduces an intermediate representation to guide text generation. A common strategy uses sparse or symbolic plans, such as structural endpoints (Brei et al., 2023), abstract narrative function sequences (Wang and Kreminski, 2024), or dynamically controlled outlines (Wang et al., 2023). While these methods improve structural control, their plans are "skeletons" with low information density, creating a substantial semantic gap between planning and execution. Some work attempts to bridge this gap through richer intermediate representations. For example, DOC (Yang et al., 2023b) proposes generating a highly detailed outline prior to story generation. However, even with enhanced detail, the intermediate representation remains an outline, which is a non-narrative, structural abstraction. Other approaches employ more sophisticated pipelines. CML-BENCH (Zheng et al., 2025) proposes a modular pipeline mirroring the human creative workflow, while frameworks like HoLLMwood (Kor et al., 2023) and Agents’ Room (Huot et al., 2024) use multi-agent architectures. Other work, Similarly, R2 (Lin et al., 2024) uses causal plot graphs as intermediate structures. Despite their architectural differences, all these methods ultimately rely on lossy, non-narrative plans to generate final prose, failing to fully bridge the semantic gap between planning and execution. An alternative paradigm focuses on iterative refinement. A representative example, Re3 (Yang et al., 2022), employs a recursive process of "reprompting and revision" instructing models to critique and improve their own outputs. This approach attempts to fix flaws after generation through complex and computationally expensive revision cycles, rather than preventing them during generation.
All these approaches, whether using sparse plans, detailed outlines, or iterative revision, highlight the persistent challenge of ensuring narrative quality in autonomous generation. Our framework addresses this challenge through architectural redesign. We leverage novel as a dense, narratively-complete intermediate state, effectively mitigating the representational gap that exists in sparse planning methods. Crucially, this design decouples two disparate skills into separate stages: the first stage handles creative narrative generation (outline-to-novel), while the second focuses exclusively on format conversion (novel-to-screenplay). Each stage thus addresses one distinct capability, transforming screenplay generation from a task requiring simultaneous mastery of conflicting constraints into a two-stage sequential process.
Data Synthesis for Narrative Tasks. The effectiveness of any generation framework depends critically on high-quality training data, which remains particularly scarce for specialized tasks like screenplay generation. To address this scarcity, researchers have developed innovative data synthesis strategies.
One research direction focuses on reverse synthesis, or reverse-engineering plans from existing texts. For example, EIPE-text (You et al., 2023) iteratively refines plans extracted from a corpus using a QA-based evaluation, while Ex3 (Huang et al., 2023) extracts hierarchical outlines from novels to construct training pairs. A notable variant is the work of Ahuja et al. (2024), which synthesizes "flawed" data to create a benchmark for plot hole detection, demonstrating the usefulness of synthetic data for complex narrative analysis tasks.
Another research direction employs forward synthesis, using LLMs to generate new training data. RLCD (Yang et al., 2023a) generates contrastive preference pairs from positive and negative prompts to support reinforcement learning, while the work of Zhu et al. (2024) uses an elaborate procedural pipeline to generate large-scale plot data for model distillation.
While effective, these synthesis techniques focus on single-stage generation and do not address the data requirements of multi-stage frameworks like ours. Our two-stage approach requires outline-to-novel training data, which existing screenplay datasets cannot provide as they contain only final screenplay format without intermediate novel representations. We resolve this through a hybrid synthesis strategy that integrates both research directions. The process begins with reverse synthesis, extracting structured inputs and narrative directives from existing screenplays. Forward synthesis then generates high-quality, novel-style narratives from these components, producing the outline-to-novel training pairs needed for our first-stage model.
In summary, existing research reveals gaps in the completeness of task decoupling, the richness of intermediate representations, and the targeted application of data synthesis. Our work addresses these three challenges through a unified framework that introduces a two-stage pipeline, leverages the novel as a dense intermediate state, and employs a hybrid data synthesis strategy tailored to support this unique architecture.
@@ 3 Methodology
This section presents our proposed methodology DSR, a framework designed to generate well-crafted screenplays. We begin by formally defining the screenwriting task and critically examining the limitations inherent in direct end-to-end generation. This analysis establishes the foundational rationale for moving beyond direct approaches and adopting a decomposed strategy. Building on this foundation, we introduce the core architecture of the DSR framework, which features a two-stage decomposed design that decouples creative narrative generation from stylistic format conversion through the use of novel-style prose as an intermediate representation. Finally, to enable practical implementation of this architecture, we present the hybrid data synthesis strategy developed to support it. This strategy encompasses both the reverse synthesis and forward synthesis processes, which are critical for generating the necessary training data.
@@ 3.2 The DSR Framework
The choice of novelistic prose as the intermediate representation is a central component of our framework’s design. This choice is motivated by two observations about LLMs (OpenAI et al., 2024; Touvron et al., 2023). First, while LLMs are pre-trained on vast corpora of narrative texts such as novels and stories, they have seen comparatively limited well-formatted screenplay data during training. Second, there exists a significant difference in information density between the two formats: novels can unfold plots with extensive descriptions and details, whereas screenplays must compress the same narrative into a much more concise format with strict structural constraints. Therefore, we introduce the novel as an intermediate representation, decomposing the generation task into two subtasks: first generating a novel from the outline, then converting the novel into a screenplay.
@@
The first stage, Stage 1: Outline-to-Novel Expansion, trains a model $M_{\theta_{1}}$ to perform both reasoning and creative generation. Its objective is to learn a mapping from the input brief $X$ to a structured output containing both a CoT analysis $I_{c}$ and the full novel prose $N$. By training the model to first generate a CoT, we encourage it to develop an internal reasoning process that enhances the quality and coherence of the subsequent novel generation. To achieve this, we train the model on a high-quality dataset $\mathcal{D}_{novel}=\{(X_{i},(I_{c,i},N_{i}))\}_{i=1}^{|\mathcal{D}_{novel}|}$. Creating such a high-quality dataset is challenging. We have access to professionally written screenplays $S$, but not the original high-level briefs $X$ used to create them. We therefore develop a hybrid data synthesis strategy illustrated in Figure 2, combining the strengths of reverse and forward synthesis to construct our training pairs. The input $X$ is created via reverse synthesis to resemble a realistic writer’s brief, while the target $N$ is created via forward synthesis to ensure high fidelity and narrative richness.
The foundation of our synthesis strategy is a high-quality, pre-processed screenplay corpus. As shown in the lower-left corner of Figure 2, all raw scripts first pass through a standardized cleaning pipeline to ensure data quality and consistency. The pipeline includes four main steps: 1) Script Selection filters for relevant screenplays, 2) Format Standardization unifies diverse writing styles into a single schema, 3) Scene Structuring parses scenes into a consistent data format, and 4) Scene Filtering removes noisy or irrelevant scenes. This resulting curated corpus contains over $200$ series with more than $50000$ scenes, serving as the source material for all subsequent synthesis tasks.
@@
Part A: Reverse Synthesis of Inputs ($X$) and Narrative Directives ($I_{c}$) As depicted in Step 1: Reverse Compression of Figure 2, our process begins by taking a professional screenplay $S$ from the pre-processed corpus and reverse-engineering the corresponding input $X$ to simulate a writer’s outline. However, reverse synthesis presents several key challenges. A primary difficulty lies in achieving the correct level of detail: the input must provide enough critical information (e.g., character conflicts and plot turning points) without including so much that the task becomes unrealistic. Moreover, reliance on LLMs (e.g., GPT-4) introduces the risk of hallucinations that create details inconsistent with the original screenplay, thereby reducing the quality of training data. This is further complicated by the "iceberg problem" of information sufficiency: a written screenplay is merely a condensation of a much larger volume of unwritten thought, and a simple summary fails to capture the essential underlying context.
@@
Upon generating the rich narrative prose $N$ from Stage 1, the second stage, Stage 2: Novel-to-Screenplay Conversion, performs the stylistic transformation. This is an inference-only stage that requires no additional fine-tuning. We leverage the powerful in-context learning capabilities of a separate large-scale model, denoted as $M_{api}$ (e.g., GPT-4), to approximate the distribution $P(S|N,X)$. The conversion is guided by a carefully engineered prompt $\pi$, which instructs the model to act as a professional screenwriter and provides clear formatting rules. The final screenplay $S$ is generated by sampling from this model:

### v2

@@ Abstract
The screenplay serves as the foundation for television production, defining narrative structure, character development, and dialogue. While Large Language Models (LLMs) offer great potential in this creative process, direct end-to-end generation approaches often fail to produce well-crafted screenplays. We argue this failure stems from forcing a single model to simultaneously master two disparate capabilities: creative narrative construction and rigid format adherence. To enable LLMs to generate high-quality screenplays, we introduce Dual-Stage Refinement (DSR), a decomposed framework that explicitly decouples creative narrative generation from format conversion. In the first stage, the framework transforms a brief outline into rich, novel-style prose. The second stage then refines this prose into a professionally formatted screenplay. This separation enables the model to specialize in one distinct capability at each stage. A significant challenge in implementing DSR is the scarcity of paired outline-to-novel data. We address this through a hybrid data synthesis strategy that combines reverse synthesis (deconstructing existing screenplays into structured inputs) and forward synthesis (generating high-quality novel-style texts as training targets). Extensive experiments show that in blind evaluations by professional screenwriters, screenplays generated by DSR achieve a $75\%$ win rate against strong baselines like Gemini-2.5-Pro and reach $82.7\%$ of human-level performance, demonstrating that decomposed generation architecture is highly effective for specializing LLMs in complex creative domains.
@@
Hang Lei1thanks: Corresponding author.   Shengyi Zong1  Zhaoyan Li1   Ziren Zhou1,2   Hao Liu1  Liang Yu1 1Alibaba Group,  2Peking University {leihang.lh,zongshengyi.zsy,lzy434483,zhouziren.zzr,lh414475,deyi.yl}@alibaba-inc.com
@@ 1 Introduction
The advent of LLMs brings new potential to screenplay generation [^0] The terms ‘screenplay’ and ‘script’ in this work refer to scripts for episodic television, not feature films.. LLMs can serve as creative assistants by handling time-consuming tasks such as generating plot variations and character biographies, allowing writers to focus on higher-level creative decisions. However, current general-purpose LLMs must be adapted to understand screenplay-specific narrative structures, maintain story coherence, and manage complex character development. The challenge is developing specialized AI tools that enable writers to produce higher-quality screenplays efficiently.
@@
A screenplay differs fundamentally from a novel as a structured framework for visual storytelling. Rather than relying on prose descriptions or internal monologue, screenplays employ "showing, not telling" through precise audiovisual language—scene headings, concise action lines, and authentic dialogue. Consequently, an effective screenplay generation model must maintain narrative coherence, ensure character consistency, and adhere to strict formatting conventions.
Current research has made notable progress in controllable story generation, developing models that follow plot outlines (rashkin2020plotmachinesoutlineconditionedgenerationdynamic) or incorporate commonsense knowledge (wang2022incorporatingcommonsenseknowledgestory). In screenwriting, systems like Dramatron (mirowski2022cowritingscreenplaystheatrescripts) employ hierarchical prompt chaining.
@@
However, these approaches focus primarily on narrative content and logical coherence. We address a more complex challenge: generating production-ready screenplays that require both creative narrative generation ("what to write") and rigid format conversion ("how to write"). Handling both skills jointly increases training complexity and often produces outputs that superficially mimic screenplay style but lack professional quality. This suggests decoupling these skills, allowing the model to focus on one task at a time.
@@
We introduce Dual Stage Refinement (DSR), a decomposed framework that separates screenplay generation into two stages. The first stage translates high-level outlines into rich, novel-style prose, focusing purely on storytelling. The second stage refines this prose into professionally formatted screenplays. This separation enables the model to master one distinct skill at each step.
@@
Implementing DSR presents a data challenge: the lack of paired outline-to-novel data for training the first stage. We address this through a hybrid data synthesis strategy combining reverse synthesis (deconstructing screenplays into structured inputs) and forward synthesis (generating high-quality novel-style targets).
@@
Blind evaluations by professional screenwriters demonstrate DSR’s effectiveness: a $75\%$ win rate against SOTA models like Gemini-2.5-Pro and Claude-Sonnet-4, reaching $82.7\%$ of human-level performance.
Our main contributions are: (1) The DSR framework that decouples screenplay generation into creative narrative development and format conversion, addressing limitations of end-to-end approaches; (2) A hybrid data synthesis strategy combining reverse and forward synthesis to resolve data scarcity; (3) Comprehensive validation showing DSR significantly outperforms strong baselines in blind evaluations by professional screenwriters.
@@ 2 Related Works
Screenplay Generation. Early screenplay generation relied on retrieval-based methods (10.1145/3507356). Modern LLM-based frameworks like Dramatron (mirowski2022cowritingscreenplaystheatrescripts) and IBSEN (han2024ibsendirectoractoragentcollaboration) enable interactive co-writing but require substantial human supervision. tian2024largelanguagemodelscapable found that while LLMs excel at surface-level qualities, they fail at "narrative intelligence" in end-to-end generation due to the difficulty of simultaneously managing creative construction and structural constraints. Unlike collaborative frameworks, our work tackles autonomous, high-quality generation.
Planning-based Narrative Generation. Plan-and-write approaches use sparse plans (brei2023returning; wang2024guiding; wang2023improving) or detailed outlines (yang2023docimprovinglongstory), but these create semantic gaps between planning and execution. Sophisticated pipelines like CML-BENCH (zheng2025cmlbenchframeworkevaluatingenhancing), HoLLMwood (kor2023hollmwood), and R2 (lin2025r) still rely on lossy, non-narrative representations. Iterative refinement methods like Re3 (yang2022re3generatinglongerstories) attempt costly post-generation fixes.
We address these limitations by using novels as dense, narratively-complete intermediate states, decoupling creative generation (outline-to-novel) from format conversion (novel-to-screenplay).
Data Synthesis for Narrative Tasks. High-quality training data is scarce for screenplay generation. Prior work employs reverse synthesis (you2023eipe; huang2023ex3; ahuja2024finding) to extract plans from texts, or forward synthesis (yang2024rlcd; zhu2024end) using LLMs to generate training data. However, these focus on single-stage generation.
Our two-stage approach requires outline-to-novel pairs unavailable in existing datasets. We employ hybrid synthesis: reverse synthesis extracts structured inputs from screenplays, then forward synthesis generates novel-style narratives, producing the training pairs needed for our framework.
@@ 3 Methodology
This section presents DSR, our proposed framework for generating well-crafted screenplays. We begin by formally defining the screenwriting task and examining the limitations of direct end-to-end generation (Section 3.1). We then introduce the DSR framework’s two-stage architecture that decouples creative narrative generation from format conversion (Section 3.2).
@@ 3.2 The DSR Framework
The choice of novelistic prose as the intermediate representation is a central component of our framework’s design. This choice is motivated by two observations about LLMs (openai2024gpt4technicalreport; touvron2023llamaopenefficientfoundation). First, while LLMs are pre-trained on vast corpora of narrative texts such as novels and stories, they have seen comparatively limited well-formatted screenplay data during training. Second, there exists a significant difference in information density between the two formats: novels can unfold plots with extensive descriptions and details, whereas screenplays must compress the same narrative into a much more concise format with strict structural constraints. Therefore, we introduce the novel as an intermediate representation, decomposing the generation task into two subtasks: first generating a novel from the outline, then converting the novel into a screenplay.
@@
The first stage Outline-to-Novel Expansion trains a model $M_{\theta_{1}}$ to perform both reasoning and creative generation. Its objective is to learn a mapping from the input brief $X$ to a structured output containing both a CoT analysis $I_{c}$ and the full novel prose $N$. By training the model to first generate a CoT, we encourage it to develop an internal reasoning process that enhances the quality and coherence of the subsequent novel generation. To achieve this, we train the model on a high-quality dataset $\mathcal{D}_{novel}=\{(X_{i},(I_{c,i},N_{i}))\}_{i=1}^{|\mathcal{D}_{novel}|}$.
Creating such a high-quality dataset is challenging. We have access to professionally written screenplays $S$, but not the original high-level briefs $X$ used to create them. We therefore develop a hybrid data synthesis strategy illustrated in Figure 2, combining the strengths of reverse and forward synthesis to construct our training pairs. The input $X$ is created via reverse synthesis to resemble a realistic writer’s brief, while the target $N$ is created via forward synthesis to ensure high fidelity and narrative richness.
The foundation of our synthesis strategy is a high-quality, pre-processed screenplay corpus. [^1] See Appendix A.1 for detailed preprocessing steps. As shown in the lower-left corner of Figure 2, all raw scripts first pass through a standardized cleaning pipeline to ensure data quality and consistency. This resulting curated corpus contains over $200$ series with more than $50000$ scenes, serving as the source material for all subsequent synthesis tasks.
@@
Part A: Reverse Synthesis of Inputs ($X$) and Narrative Directives ($I_{c}$) As depicted in Step 1: Reverse Compression of Figure 2, our process begins by taking a professional screenplay $S$ from the pre-processed corpus and reverse-engineering the corresponding input $X$ to simulate a writer’s outline. However, reverse synthesis presents several key challenges. A primary difficulty lies in achieving the correct level of detail: the input must provide enough critical information (e.g., character conflicts and plot turning points) without including so much that the task becomes unrealistic. Moreover, reliance on LLMs (e.g., GPT-4) introduces the risk of hallucinations that create details inconsistent with the original screenplay, thereby reducing the quality of training data.
@@
Upon generating the rich narrative prose $N$ from Stage 1, the second stage, Novel-to-Screenplay Conversion, performs the stylistic transformation. This is an inference-only stage that requires no additional fine-tuning. We leverage the powerful in-context learning capabilities of a separate large-scale model, denoted as $M_{api}$ (e.g., GPT-4), to approximate the distribution $P(S|N,X)$. The conversion is guided by a carefully engineered prompt $\pi$, which instructs the model to act as a professional screenwriter and provides clear formatting rules. The final screenplay $S$ is generated by sampling from this model: